First principle of differentiation
1
Expert's answer
2019-07-05T10:48:47-0400

We know that the straight line has a constant gradient or we can also say that the rate of change of y with respect to x is a constant. One can define the instantaneous rate of change of y with respect to x. This is equivalent to the following:


"\frac{{dy}}{{dx}} = \mathop {\lim }\limits_{\Delta x \to 0} \frac{{\Delta y}}{{\Delta x}}\\

\frac{{dy}}{{dx}} = \mathop {\lim }\limits_{\Delta x \to 0} \frac{{f\left( {x + \Delta x} \right) - f\left( x \right)}}{{\Delta x}} ."

It is also known as the delta method.